# The XMAGNET exascale MHD simulations of SMBH feedback in galaxy groups and clusters: Overview and preliminary cluster results

### v1

@@ 1 Introduction
The intracluster medium (ICM) – the hot, diffuse, magnetized plasma that comprises the bulk of the baryons in galaxy clusters – radiates copiously in X-ray radiation, with X-ray luminosities of massive clusters typically being in the range of $\sim 10^{43.5}-10^{45.5}$ erg/s (Fabian, 1994). “Cool core” galaxy clusters, having dense and (relatively) cool cores, are particularly X-ray luminous, with central cooling times of tens to hundreds of millions of years – far shorter than the age of the universe and the lifetime of these structures. (Hudson et al., 2010)
@@
In cool-core clusters, local thermal stability of the intracluster medium becomes a critical physical phenomenon that governs their evolution (Field, 1965). Cold gas is expected to precipitate out of the thermally unstable ICM (Voit et al., 2008; McCourt et al., 2012) and to be available to star formation and accretion onto the central SMBH (Oosterloo et al., 2024; Guo et al., 2024). An apparent consequence is that this rapidly cooling gas should result in many hundreds to thousands of Solar masses per year of cold, dense gas being available for star formation, and thus that the brightest central galaxies in clusters should see similarly high star formation rates. However, the observed star formation rate in these galaxies is lower – typically only a handful of solar masses per year (Fabian et al., 1984; Edge, 2001; Peterson et al., 2003; O’Dea et al., 2008).
@@
This observation implies that there is a heating source in the center of every cool-core galaxy cluster that offsets this cooling, and does so in a tightly-coupled way that inhibits significant amounts of gas from forming (Pizzolato & Soker, 2005). While many mechanisms have been proposed, including thermal conduction, supernova feedback, sound waves, and cluster mergers, the only mechanism that robustly fits the necessary criteria of (1) appropriate level of energy production, (2) location that would tend to deposit heat in the areas of greatest X-ray emission/cooling, and (3) ability to couple to the central cluster plasma on short time scales, is AGN feedback (Churazov et al., 2001; Bîrzan et al., 2004; Rafferty et al., 2006, 2008; Cavagnolo et al., 2008, 2010; Sun et al., 2009). This idea has been supported by observational evidence from radio surveys showing that radio-loud AGN at the centers of massive cool-core galaxies are always active and, separately, analysis of cavities created by AGN jets interacting with the ICM demonstrates that the work done by those “bubbles” over time is roughly what is needed to offset cooling in clusters (e.g., Bîrzan et al., 2004; Panagoulia et al., 2014; Prunier et al., 2024).
@@
These observations are supported by theory and simulations that have shown AGN feedback is crucial to the self-regulation of clusters (e.g., Gaspari et al., 2012, 2013; Li et al., 2015; Prasad et al., 2015; Voit et al., 2017; Meece et al., 2017; Donahue & Voit, 2022; Nobels et al., 2022; Ehlert et al., 2023). Idealized simulations of AGN feedback in galaxy clusters have demonstrated that cold gas-triggered AGN hosted by the central cluster galaxy are effective mechanisms for offsetting radiative losses in the intracluster medium, and (depending on the exact prescription for AGN triggering and feedback) result in clusters with thermodynamic states that are comparable to observed cool-core clusters (e.g., Li & Bryan, 2014; Qiu et al., 2019; Beckmann et al., 2019; Wang et al., 2021; Ehlert et al., 2023; Fournier et al., 2024). Furthermore, theory has shown that AGN feedback is important in environments less extreme than galaxy clusters. Cosmological simulations of galaxy formation show that AGN feedback is necessary to correctly reproduce the low star formation rates in massive galaxies (e.g., Naab & Ostriker, 2017; Donahue & Voit, 2022) as well as the budget of cold-phase gas (e.g., Olivares et al., 2019). It is challenging to make detailed inferences from these simulations, however, because implementations of AGN feedback operate at the resolution limit of these calculations (i.e., on $\sim$ kpc scales) and thus cannot accurately capture the interaction between the jet and the cooling, multi-phase gas. In addition, these simulations do not reproduce the morphology of the radio jets that dominate the radio sky (e.g., Shimwell et al., 2022).
@@
It is clear that there is need for a detailed examination of the impact of AGN feedback in self-regulating cool-core clusters. Multi-wavelength observations of these systems have recently shed light on the multiphase nature of the intracluster medium, showing significant complexity at relatively small spatial scales (e.g., Salomé et al., 2006; McNamara et al., 2014; Tremblay et al., 2018; Olivares et al., 2019; Vantyghem et al., 2021; Gingras et al., 2024). Filament-like structures of magnetically-supported warm ionized and cold molecular gas are expected to populate the inner tens of kpc of potentially all cool–core clusters (Olivares et al., 2019). The measured velocity structure functions and X-ray surface brightness fluctuations also imply that the ICM is turbulent, although this turbulence is subsonic (Li et al., 2020; Ganguly et al., 2023; de Vries et al., 2023).
@@
In this paper we focus on a subset of our calculations – a suite of idealized galaxy clusters with masses and other properties comparable to the Perseus cool-core cluster, with AGN feedback triggered by the formation of cool gas in the vicinity of the central supermassive black hole – and present key results from those calculations. Following papers will delve deeper into the evolution of magnetized turbulence in these clusters, closer comparisons to observation, the formation and properties of cold, magnetized filaments in the center of the clusters, and of halos with a range of different virial masses and properties.
@@ 2.2.6 AGN Feedback
where $\dot{E}_{\rm AGN}$ is the total AGN feedback rate; $\dot{E}_{T}$, $\dot{E}_{K}$, and $\dot{E}_{B}$ are the total thermal, kinetic, and magnetic AGN feedback rates; and $f_{T}$, $f_{K}$, and $f_{B}$ are the thermal, kinetic, and magnetic fractions of the total AGN feedback rate. Given the scales in our simulations we set $f_{T}=0.25$, $f_{K}=0.74$ and $f_{B}=0.01$ (or $f_{K}=0.75$ and $f_{B}=0$ in the hydrodynamic case).
@@ 2.2.7 Thermal AGN Feedback
In the thermal feedback model, thermal energy is deposited volumetrically within a sphere around the center of the halo where the presumed AGN resides.
@@ 2.2.8 Kinetic AGN Feedback
In the kinetic feedback model, kinetic energy and mass is injected into two disks offset above and below the midplane of the AGN jet. These disks have a radius of $R_{K}$, each with thickness $H_{K}$, and are offset from the midplane by $h_{0,K}$ such that kinetic feedback is injected as far away from the midplane as $H_{K}+h_{0,K}$. The rate of mass-energy injected by the jet is set proportional the fraction of accreted mass afforded to kinetic feedback, with $\epsilon$ fraction of that mass-energy proportioned to energy, i,.e.,
@@
where $c$ is the speed of light. The jet then injects a mass density
@@
where $R_{K}=500$ pc and $H_{K}=2\Delta_{x}\approx 200$ pc. The jet speed of the injected material is given by:
@@ 2.2.9 Magnetic AGN Feedback
to be divergence-free to machine precision. For all simulations the scales are chosen444 A more complex pinching magnetic tower model, as well as mass injection associated with the magnetic feedback channel, are also implemented in the code but not used in the simulations presented in this paper given their large-scale focus. such that the magnetic donut is effectively seeded within the kinetic jet launching region, i.e., $h_{\mathrm{0,M}}=h_{\mathrm{0,K}}=1$ kpc, $h_{\mathrm{M}}=h_{\mathrm{K}}=2\Delta_{x}\approx 200$ pc, and $L_{M}=250$ pc.
@@ 2.2.10 AGN cold mass triggering
where $\epsilon_{\rm AGN}=10^{-3}$ is the efficiency with which rest-mass from the accreted cold gas is turned into feedback energy.
@@
For model completeness we implemented other types of accretion, e.g., “boosted Bondi,” but thsese are not used in the present simulations.
@@ 2.2.11 Stellar feedback
Heat and mass feedback from Type Ia supernovae (SNIa) are injected in a spherically symmetric kernel proportional to the (fixed) stellar mass density of the central cluster galaxy. The energy density and mass density injected into a cell at distance $r$ from the cluster center is
@@ 2.2.12 Under the rug: on floors and ceilings
Given the extreme conditions in the central region of the simulation, we limit certain quantities to remain within the reasonable limits of the overall equation system solved (e.g., not entering a relativistic regime). More specifically, within $r<20$ kpc, the velocity is limited to $0.05c$, the Alfvén velocity is also limited to $0.05c$, and the temperature limited to $5\times 10^{9}$ K, which are implemented as ceilings following the feedback routines. To measure the impact of these limits we tracked the amount of energy being removed from the simulation. Their effective powers were signifcantly smaller ($\ll 10^{3}$) than the AGN feedback and, thus, negligible for the dynamics in the simulation. In addition, we employ a global temperature floor at $10^{4}$ K, which is below the lower end of the cooling tables.
@@ 2.3 The simulation suite
A systematic overview of the initial conditions of all current simulations in the XMAGNET suite are given in Table 1. All the cluster simulations, i.e., Perseus-like, a larger one ($\mathtt{Lg}$), and a smaller one ($\mathtt{Sm}$), use a uniform $2{,}560^{3}$ mesh covering the central region whereas in the group simulations, $\mathtt{SPG}$ and $\mathtt{MPG}$, only the $[50\,\mathrm{kpc}]^{3}$ are covered by a $512^{3}$ mesh at the highest resolution of $\Delta_{x}\approx 100$ pc in the center. Similarly, the initial velocity perturbations in the group simulation have been scaled down to 30 km/s at 40 kpc scales.
@@
Sim. $B_{0}$ $Z$ $K_{0}$ $K_{100}$ $\alpha_{K}$ $r_{\mathrm{200}}$ $M_{\mathrm{NFW}}$ $c_{\mathrm{NFW}}$ $M_{\mathrm{BCG}}$ $r_{\mathrm{BCG}}$ $M_{\mathrm{SMBH}}$
@@
Table 1: Key differentiating parameters of all simulations: The initial magnetic field strength is the RMS value of $B_{0}$ on 100 kpc scales, $Z$ is the metallicity of the cooling table, $K_{0}$ and $K_{100}$ the specific entropies in the center and at $r=100$ kpc, respectively, $\alpha_{K}$ is slope of the entropy profile, $r_{\mathrm{200}}$ is the virial radius where the initial density of $\rho(r_{\mathrm{ref}})=9.47\times 10^{-29}$ g/cm3 is fixed, $M_{\mathrm{NFW}}(=M_{200}$), $M_{\mathrm{BCG}}$, and $M_{\mathrm{SMBH}}$ are the dark matter halo, BCG, and SMBH masses, $c_{\mathrm{NFW}}$ is the concentration parameter, and $r_{\mathrm{BCG}}$ is the scale radius of the Hernquist profile for the BCG.
@@ 3.2 Thermodynamics and energetics
The overall energy balance is depicted in Fig. 8 based on the median radial profiles of the sonic Mach number (as a proxy for the kinetic energy over thermal pressure), the Alfvénic Mach (as proxy for the kinetic over magnetic energy), and plasma beta (as thermal pressure over magnetic energy/pressure). The Alvénic Mach is comparable within $r\leq 4$ kpc (dropping from several tens to several few) for all simulations that include magnetic fields. Then it continues to decline, eventually becoming sub-Alfvénic beyond $r\gtrsim 100$ kpc in the Fiducial and WeakCool case whereas for WeakMag it increases again rapidly to $\sim 200$, i.e., to a highly super-Alfvénic state. This is also reflected in the plasma beta. For WeakCool it remains approximately constant within $r\leq 500$ kpc with values of several hundreds up to 1000. At $r>10$ kpc this also applies to the Fiducial case; however, at small distances from the center the plasma beta drops to $\mathcal{O}(10)$. Again, the WeakMag simulation stands out with plasma betas reaching $10^{5}$ beyond $r>10$ kpc. Given the $1/B^{2}$ scaling of the plasma beta and an initial difference in $B_{0}$ of $10^{3}$ between the Fiducial/WeakCool and WeakMag case, this already gives a first indication of an effective magnetic field amplification that is further discussed in Subsection 3.4.

### v2

@@ 1 Introduction
The intracluster medium (ICM) – the hot, diffuse, magnetized plasma that comprises the bulk of the baryons in galaxy clusters – radiates copiously in X-ray radiation, with X-ray luminosities of massive clusters typically being in the range of $\sim 10^{43.5}-10^{45.5}$ erg/s (Fabian, 1994). “Cool core” galaxy clusters, having dense and (relatively) cool cores, are particularly X-ray luminous, with central cooling times of tens to hundreds of millions of years – far shorter than the age of the universe and the lifetime of these structures (Hudson et al., 2010; Cavagnolo et al., 2008).
@@
In cool-core clusters, local thermal stability of the intracluster medium becomes a critical physical phenomenon that governs their evolution (Field, 1965). Cold gas is expected to precipitate out of the thermally unstable ICM (Voit et al., 2008; McCourt et al., 2012; Sharma et al., 2012) and to be available to star formation and accretion onto the central SMBH (Soker et al., 2001; Oosterloo et al., 2024; Guo et al., 2024). An apparent consequence is that the rapidly cooling gas would result in many hundreds to thousands of solar masses per year of cold, dense gas being available for star formation. This would result in brightest cluster galaxies having a very high star formation rate ($\sim$ several $100s$ M yr-1), and thus it was predicted that the brightest central galaxies in clusters should see similarly high star formation rates. The observed star formation rate in these galaxies is lower, however – typically only a handful of solar masses per year (Fabian et al., 1984; Edge, 2001; Peterson et al., 2003; O’Dea et al., 2008).
@@
This observation implies that there is a heating source in the center of every cool-core galaxy cluster that offsets this cooling, and does so in a tightly-coupled way that inhibits significant amounts of gas from forming (Pizzolato & Soker, 2005). Although several mechanisms have been proposed, including thermal conduction (Voigt & Fabian, 2004; Ruszkowski & Oh, 2011), supernova feedback (Voit & Bryan, 2001), sound waves (Ruszkowski et al., 2004), infalling galaxies and sub-halos (Dekel & Birnboim, 2008), and cluster mergers, the only mechanism that robustly fits the necessary criteria of (1) appropriate level of energy production, (2) location that would tend to deposit heat in the areas of greatest X-ray emission/cooling, and (3) ability to couple to the central cluster plasma on short time scales, is AGN feedback (Churazov et al., 2001; Bîrzan et al., 2004; Rafferty et al., 2006, 2008; Cavagnolo et al., 2008, 2010; Sun et al., 2009). This idea has been supported by observational evidence from radio surveys showing that radio-loud AGN at the centers of massive cool-core galaxies are always active and, separately, analysis of cavities created by AGN jets interacting with the ICM demonstrates that the work done by those “bubbles” over time is roughly what is needed to offset cooling in clusters (e.g., Bîrzan et al., 2004; Panagoulia et al., 2014; Prunier et al., 2024).
@@
These observations are supported by theory and simulations that have shown AGN feedback is crucial to the self-regulation of clusters (e.g., Gaspari et al., 2012; Li et al., 2015; Prasad et al., 2015; Yang & Reynolds, 2016; Meece et al., 2017; Chadayammuri et al., 2021; Nobels et al., 2022; Ehlert et al., 2023). Idealized simulations of AGN feedback in galaxy clusters have demonstrated that cold gas-triggered AGN hosted by the central cluster galaxy are effective mechanisms for offsetting radiative losses in the intracluster medium, and (depending on the exact prescription for AGN triggering and feedback) result in clusters with thermodynamic states that are comparable to observed cool-core clusters (e.g., Li & Bryan, 2014; Qiu et al., 2019; Beckmann et al., 2019; Wang et al., 2021; Ehlert et al., 2023; Fournier et al., 2024). Furthermore, theory has shown that AGN feedback is important in environments less extreme than galaxy clusters. Cosmological simulations of galaxy formation show that AGN feedback is necessary to correctly reproduce the low star formation rates in massive galaxies (e.g., Naab & Ostriker, 2017; Weinberger et al., 2018; Donahue & Voit, 2022; Piotrowska et al., 2022) as well as the budget of cold-phase gas (e.g., Olivares et al., 2019). It is challenging to make detailed inferences from these simulations, however, because implementations of AGN feedback operate at the resolution limit of these calculations (i.e., on $\sim$ kpc scales) and thus cannot accurately capture the interaction between the jet and the cooling, multi-phase gas. In addition, these simulations do not reproduce the morphology of the radio jets that dominate the radio sky (e.g., Shimwell et al., 2022).
@@
It is clear that there is need for a detailed examination of the impact of AGN feedback in self-regulating cool-core clusters. Multi-wavelength observations of these systems have recently shed light on the multiphase nature of the intracluster medium, showing significant complexity at relatively small spatial scales (e.g., Salomé et al., 2006; McNamara et al., 2014; Tremblay et al., 2018; Olivares et al., 2019; Vantyghem et al., 2021; Gingras et al., 2024). Filament-like structures of magnetically-supported warm ionized and cold molecular gas are expected to populate the inner tens of kpc of potentially all cool–core clusters (Russell et al., 2014; Olivares et al., 2019), with cold molecular gas also sometimes observed in much smaller, disk-like structures (see, e.g., Hamer et al., 2014). The measured velocity structure functions and X-ray surface brightness fluctuations also imply that the ICM is turbulent, although this turbulence is subsonic (Li et al., 2020; Ganguly et al., 2023; de Vries et al., 2023).
@@
In this paper we focus on a subset of our calculations – a suite of idealized galaxy clusters with masses and other properties comparable to the Perseus cool-core cluster, with AGN feedback triggered by the formation of cool gas in the vicinity of the central supermassive black hole – and present key results from those calculations. Following papers will delve deeper into the the velocity structure functions of the hot and cold phases and projection effects (Fournier et al., 2025), evolution of magnetized turbulence in these clusters, closer comparisons to observation, the formation and properties of cold, magnetized filaments in the center of the clusters, and of halos with a range of different virial masses and properties.
@@ 2.2.6 AGN Feedback
where $\dot{E}_{\rm AGN}$ is the total AGN feedback rate; $\dot{E}_{T}$, $\dot{E}_{K}$, and $\dot{E}_{B}$ are the total thermal, kinetic, and magnetic AGN feedback rates; and $f_{T}$, $f_{K}$, and $f_{B}$ are the thermal, kinetic, and magnetic fractions of the total AGN feedback rate. Given the scales in our simulations we set $f_{T}=0.25$, $f_{K}=0.74$ and $f_{B}=0.01$ (or $f_{K}=0.75$ and $f_{B}=0$ in the hydrodynamic case). The thermal and kinetic fractions are motivated by Meece et al. (2017) and the magnetic fraction was chosen to be subdominant given the distance to the SMBH’s accretion disk. The value of $f_{B}$ is in agreement with general relativistic magnetohydrodynamical simulations of black hole-launched jets (see, e.g., Liska et al., 2020; Kaaz et al., 2023), where the jet is magnetically-driven at a few gravitational radii but is almost entirely kinetic at tens of thousands of gravitational radii.
@@ 2.2.7 Thermal AGN Feedback
In the thermal feedback model, thermal energy is deposited volumetrically within a sphere around the center of the halo where the presumed AGN resides (Booth & Schaye, 2009).
@@ 2.2.8 Kinetic AGN Feedback
In the kinetic feedback model, kinetic energy and mass is injected into two disks offset above and below the midplane of the AGN jet. These disks have a radius of $R_{K}$, each with thickness $H_{K}$, and are offset from the midplane by $h_{0,K}$ such that kinetic feedback is injected as far away from the midplane as $H_{K}+h_{0,K}$. Following Meece et al. (2017), the rate of mass-energy injected by the jet is set proportional the fraction of accreted mass afforded to kinetic feedback, with $\epsilon$ fraction of that mass-energy proportioned to energy, i,.e.,
@@
where $c$ is the speed of light. The jet then injects a mass density at the rate of
@@
where $R_{K}=500$ pc and $H_{K}=2\Delta_{x}\approx 200$ pc – chosen so that the jet and associated physical instabilities at the boundary layer are well resolved , cf., Weinberger et al. (2023). The jet speed of the injected material is given by:
@@ 2.2.9 Magnetic AGN Feedback
to be divergence-free to machine precision. For all simulations the scales are chosen444 A more complex pinching magnetic tower model, as well as mass injection associated with the magnetic feedback channel, are also implemented in the code but not used in the simulations presented in this paper given their large-scale focus. such that the magnetic donut is effectively seeded within the kinetic jet launching region, i.e., $h_{\mathrm{0,M}}=h_{\mathrm{0,K}}=1$ kpc, $h_{\mathrm{M}}=h_{\mathrm{K}}=2\Delta_{x}\approx 200$ pc, and $L_{M}=250$ pc, similar to Weinberger et al. (2017).
@@ 2.2.10 AGN cold mass triggering
where $\epsilon_{\rm AGN}=10^{-3}$ is the efficiency with which rest-mass from the accreted cold gas is turned into feedback energy following, e.g., Gaspari et al. (2013); Li & Bryan (2014); Meece et al. (2017); Wang et al. (2021).
@@
For model completeness we implemented other types of accretion, e.g., “boosted Bondi,” but these are not used in the present simulations.
@@ 2.2.11 Stellar feedback
Heat and mass feedback from Type Ia supernovae (SNIa) are injected in a spherically symmetric kernel proportional to the (fixed) stellar mass density of the central cluster galaxy. Following Voit et al. (2015); Prasad et al. (2020), the energy density and mass density injected into a cell at distance $r$ from the cluster center is
@@ 2.2.12 Under the rug: on floors and ceilings
Given the extreme conditions in the central region of the simulation, we limit certain quantities to remain within the reasonable limits of the overall equation system solved (e.g., not entering a relativistic regime). More specifically, within $r<20$ kpc, the velocity is limited to $0.05c$, the Alfvén velocity is also limited to $0.05c$, and the temperature limited to $5\times 10^{9}$ K, which are implemented as ceilings following the feedback routines. To measure the impact of these limits we tracked the amount of energy being removed from the simulation. Their effective powers were significantly smaller ($\ll 10^{3}$) than the AGN feedback and, thus, negligible for the dynamics in the simulation. In addition, we employ a global temperature floor at $10^{4}$ K, which is below the lower end of the cooling tables.
@@ 2.3 The simulation suite
A systematic overview of the initial conditions of all current simulations in the XMAGNET suite are given in Table 1. All the cluster simulations, i.e., Perseus-like, a larger one ($\mathtt{Lg}$), and a smaller one ($\mathtt{Sm}$), use a uniform $2{,}560^{3}$ mesh covering the central region, see Fig. 2 for an illustration, whereas in the group simulations, $\mathtt{SPG}$ and $\mathtt{MPG}$, only the $[50\,\mathrm{kpc}]^{3}$ are covered by a $512^{3}$ mesh at the highest resolution of $\Delta_{x}\approx 100$ pc in the center. Similarly, the initial velocity perturbations in the group simulation have been scaled down to 30 km/s at 40 kpc scales.
@@
Sim. $B_{0}$ $Z$ $K_{0}$ $K_{100}$ $\alpha_{K}$ $r_{\mathrm{ref}}$ $M_{\mathrm{NFW}}$ $c_{\mathrm{NFW}}$ $M_{\mathrm{BCG}}$ $r_{\mathrm{BCG}}$ $M_{\mathrm{SMBH}}$
@@
Table 1: Key differentiating parameters of all simulations: The initial magnetic field strength is the RMS value of $B_{0}$ on 100 kpc scales, $Z$ is the metallicity of the cooling table, $K_{0}$ and $K_{100}$ the specific entropies in the center and at $r=100$ kpc, respectively, $\alpha_{K}$ is slope of the entropy profile, $r_{\mathrm{ref}}$ is the radius where the initial density of $\rho(r_{\mathrm{ref}})=9.47\times 10^{-29}$ g/cm3 is fixed, $M_{\mathrm{NFW}}(=M_{200}$), $M_{\mathrm{BCG}}$, and $M_{\mathrm{SMBH}}$ are the dark matter halo, BCG, and SMBH masses, $c_{\mathrm{NFW}}$ is the concentration parameter, and $r_{\mathrm{BCG}}$ is the scale radius of the Hernquist profile for the BCG.
@@ 3.2 Thermodynamics and energetics
The overall energy balance is depicted in Fig. 8 based on the median radial profiles of the sonic Mach number (as a proxy for the kinetic energy over thermal pressure), the Alfvénic Mach (as proxy for the kinetic over magnetic energy), and plasma beta (as thermal pressure over magnetic energy/pressure). The Alvénic Mach is comparable within $r\leq 4$ kpc (dropping from several tens to several few) for all simulations that include magnetic fields. Then it continues to decline, eventually becoming sub-Alfvénic beyond $r\gtrsim 100$ kpc in the Fiducial and WeakCool case whereas for WeakMag it increases again rapidly to $\sim 200$, i.e., to a highly super-Alfvénic state. This is also reflected in the plasma beta. For WeakCool it remains approximately constant within $r\leq 500$ kpc with values of several hundreds up to 1000. At $r>10$ kpc this also applies to the Fiducial case. However, at small distances from the center the plasma beta drops to $\mathcal{O}(10)$ in the Fiducial case and $\mathcal{O}(100)$ in the WeakCool case, which is tied to the larger amont of cold gas in the former compared to the latter. Again, the WeakMag simulation stands out with plasma betas reaching $10^{5}$ beyond $r>10$ kpc. Given the $1/B^{2}$ scaling of the plasma beta and an initial difference in $B_{0}$ of $10^{3}$ between the Fiducial/WeakCool and WeakMag case, this already gives a first indication of an effective magnetic field amplification that is further discussed in Subsection 3.4.